4. Suppose Σί, 1 Xi-1, ΣΙ-, x-2, and n-5. Evaluate the followings: a) Σί-1[Xi + X) b) 5. Let X-1 Σ., x1-1 with n-100. (i) Obtain ΣΙ.i Xu (ii) Evaluate Σ.1x,-X) and X +1
Answer #1

4/a).

∑Xi=1, ∑Yi=2 and “n=5”, => ∑(Xi+Yi) = ∑Xi + ∑Yi = 1+2 = 3, => ∑(Xi + Yi) = 3.

b).

∑(Xi-2*Yi) = ∑Xi - ∑2*Yi = ∑Xi – 2*∑Yi = 1 – 2*2 = (-3), => ∑(Xi - 2*Yi) = (-3).

c).

∑(Xi - 1), => ∑Xi - ∑1 = 1 – n = 1-5 = (-4), => ∑(Xi - 1) = (-4).

5/i).

=> X bar = (1/n)*∑Xi = 1, => ∑Xi = 1*n = 100, => ∑Xi = 100.

ii).

=> ∑(Xi – X bar) = ∑Xi – ∑X bar = ∑Xi – X bar*∑1 = 100 – 1*n = 100 – 1*100 = 0.

=> ∑(Xi – X bar) = 0.

and

=> ∑(X bar + 1) = ∑X bar + ∑1 = X bar*∑1 + n = 1*n + n = 2n = 200, => ∑(X bar + 1) = 200.